OpenStudy (anonymous):

I cant take it anymore. Can someone please walk me through on how to simplify 8 over 2 plus 2i.

OpenStudy (anonymous):

Multiply both the top and bottom by 2-2i so they cancel out

OpenStudy (anonymous):

By the way, if you hear the word "conjugate" in class, what it means is to change the sign in the middle of a binomial or complex number. 2-2i is the conjugate of 2+2i.

jimthompson5910 (jim_thompson5910):

Multiply top and bottom by 2-2i to get... \[\Large \frac{8}{2+2i}\] \[\Large \frac{8(2-2i)}{(2+2i)(2-2i)}\] \[\Large \frac{8(2-2i)}{(2)^2-(2i)^2}\] \[\Large \frac{8(2-2i)}{4-4i^2}\] \[\Large \frac{8(2-2i)}{4-4(-1)}\] \[\Large \frac{8(2-2i)}{4+4}\] \[\Large \frac{8(2-2i)}{8}\] \[\Large \frac{\cancel{8}(2-2i)}{\cancel{8}}\] \[\Large 2-2i\] Therefore, \[\Large \frac{8}{2+2i} = 2-2i\] Notes: * We multiply top and bottom by the denominators conjugate. The conjugate of a+bi is a-bi * In step 3, I'm using the difference of squares rule (a+b)(a-b) = a^2 - b^2 * \(\Large i = \sqrt{-1}\), so \(\Large i^2 = -1\)

OpenStudy (anonymous):

oh wow. thank you very much for showing me the process.i can use this to do other problems as well. thank you

jimthompson5910 (jim_thompson5910):

you're welcome